OpenStudy (anonymous):

evaluate log 16, given that log 2=0.4307 u u how is this done?

OpenStudy (anonymous):

Little bit long but try this way- Let log u (16)=x Then 16=u^x 2^4=u^x so, 2=u^(x/4) Now log (u) 2=0.4307 gives 2=u^0.4307 so u^(x/4)=u^0.4307 so x/4=0.4307 x=1.7228

OpenStudy (anonymous):

16 = 2^4-> 4 log 2...

OpenStudy (anonymous):

Hence the required value of the logarithm is 1.7228

OpenStudy (anonymous):

thank you....i understood

OpenStudy (anonymous):

Good, kinda messy though!

OpenStudy (anonymous):

Oh, I see the u's are part of the question...duh. They are just floating alone in my browser.

myininaya (myininaya):

\[\log_u(16)=\log_u(2^4)=4\log_u(2)=4[0.4307]=1.7228\]

OpenStudy (anonymous):

So I was right, it's a one liner...

myininaya (myininaya):

lol